Grade 11General Physics

A particle moves in a straight lineand its position is given as x^2=1+p^2 ,prove that acceleration of particle is 1/x^3

Annu
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
x² = 1 + p²
x = sqrt (1 +p²)
v = dx/dt = 2p/sqrt(1 + p²)
a = dv/dt = d²x/dt² = [2 sqrt(1 +p²) - 4p² /sqrt(1+p²) ]/1 +p²
= 2 - 2p² / (1+p²)^3/2
acceleration is proportional to 1/x³